### v2

# Enhancing Adversarial Transferability via Component-Wise Transformation

Abstract
Deep Neural Networks (DNNs) are highly vulnerable to adversarial examples, which pose significant challenges in security-sensitive applications. Among various adversarial attack strategies, input transformation-based attacks have demonstrated remarkable effectiveness in enhancing adversarial transferability. However, existing methods still perform poorly across different architectures, even though they have achieved promising results within the same architecture. This limitation arises because, while models of the same architecture may focus on different regions of the object, the variation is even more pronounced across different architectures. Unfortunately, current approaches fail to effectively guide models to attend to these diverse regions. To address this issue, this paper proposes a novel input transformation-based attack method, termed Component-Wise Transformation (CWT). CWT applies interpolation and selective rotation to individual image blocks, ensuring that each transformed image highlights different target regions, thereby improving the transferability of adversarial examples. Extensive experiments on the standard ImageNet dataset show that CWT consistently outperforms state-of-the-art methods in both attack success rates and stability across CNN- and Transformer-based models.
1 Introduction
With the rapid development of deep learning, artificial intelligence has achieved significant progress in computer vision, finding widespread applications in tasks such as image classification [12, 14], object detection [31, 30], and semantic segmentation [32, 23]. However, the adversarial vulnerability of deep learning models has gradually surfaced as a critical issue, limiting their deployment in real-world scenarios. Research shows that carefully crafted adversarial examples—created by adding imperceptible perturbations to normal input data—can cause deep learning models to make erroneous predictions [37, 10, 11]. This phenomenon not only affects model performance on standard test datasets but poses severe risks in high-stakes applications [6, 53, 36] such as autonomous driving and medical diagnostics.
Figure 1: There shows our CWT process for image transformation. The heatmaps generated on ResNet-18. Our method successfully enables a single surrogate model to focus on different regions of an object.
In black-box scenarios, attackers lack access to the details of the target models and can only rely on a limited number of surrogate models to craft adversarial examples. Therefore, a critical challenge is transferability, which refers to the ability of adversarial examples generated on surrogate models to deceive other target models, even when the target models have different architectures or are trained on different datasets [50, 41].
To improve adversarial transferability, researchers have proposed a variety of methods for generating adversarial examples. Gradient-based methods [10, 17, 4, 20, 8] leverage the gradient information of models to craft adversarial examples with high computational efficiency and serve as the foundation for many other approaches. However, they generally suffer from poor transferability, particularly across models with different architectures. Model-related methods [48, 47, 24] exploit internal model features, such as skip connections or backpropagation, to generate architecture-specific adversarial examples. While these methods perform well for specific architectures, they often overfit and lack generalizability. Ensemble-based [21, 52, 18, 1] and generation-based [33] methods aggregate the outputs of multiple models or use generative models like GANs to directly generate adversarial examples. While these approaches can improve transferability, they are constrained by high computational costs. In contrast, input transformation-based methods [50, 50, 20, 54] enhance cross-model transferability through enriching adversarial example diversity with operations such as cropping, scaling, and rotation. These methods are computationally efficient and do not require access to model parameters, making them particularly suitable for black-box attack scenarios.
Figure 2: The heatmaps generated on ResNet-18 of raw image and its transformed images by DIM, SIM, Adimx, MaskBlock, US-MM, BSR, and our proposed CWT. Notably, CWT successfully shifts the surrogate model’s attention toward the central region of the object.
However, existing input transformation-based methods still face critical challenges. As illustrated in Figure 2, some approaches fail to effectively alter the model’s attention, resulting in an attention distribution that remains consistent with the original image. In contrast, others misdirect attention to irrelevant regions of the object, deviating from the areas of interest in the original image.
To address these issues, we propose a novel method called Component-Wise Transformation (CWT), which enhances the transferability of adversarial examples by encouraging models to focus on diverse regions of the object in the original image. Specifically, CWT applies block-wise transformations that interpolate and selectively rotate image patches while preserving essential semantic information. By introducing these localized transformations, our method generates adversarial examples with enriched attention distributions that generalize effectively across different model architectures.
Our contributions can be summarized as follows:
  1. 1.
    We introduce a novel perspective on addressing the challenge of untargeted adversarial transferability. Instead of broadly enriching input images, our approach focuses on generating transformed images that enable a single surrogate model to attend to different regions of the object across various transformed images.
  2. 2.
    We propose CWT, an innovative method combining interpolation and selective rotation to enhance adversarial transferability by diversifying attention distributions.
  3. 3.
    Our method achieves state-of-the-art results on the ImageNet dataset, demonstrating superior attack success rates and lower standard deviations compared to existing approaches.
2 Related Work
Figure 3: Demonstration of the different discriminative regions of the different models. We adopt Grad-CAM [34] to visualize the attention maps of four CNN-based models—ResNet-18, ResNext-50, ResNet-101, DenseNet-121 and two Transformer-base models-Swin,Vit.
2.1 Input Transformation-based Attacks
Input transformation-based methods have been widely studied to enhance adversarial transferability, particularly in black-box settings. These methods diversify the input samples used for gradient computation, enabling adversarial examples to generalize across different model architectures. We classify these methods into four categories: scale-based methods, mixed image-based methods, block-based methods, and adaptive-based methods. Scale-based methods [50, 20] focus on resizing and transforming input images to introduce diversity in gradient computation. Mixed image-based methods [42, 40, 54] enrich input diversity by incorporating features from other images. However, whole-image transformations alone may not effectively alter the model’s attention. To address this, Block-based methods [7, 39] focus on partitioning input images into blocks and applying transformations to each block independently. Furthermore, to better utilize existing transformation methods, adaptive-based methods [46, 44, 56] introduce learnable transformation strategies to optimize adversarial transferability.
2.2 Adversarial Defense
Various defense approaches have been proposed to mitigate the threat of adversarial attacks, including adversarial training [25, 43], feature denoising [49], certified defenses [29, 11] and diffusion purification [45]. Among these, AT [35] is one of the most effective methods, where adversarial examples are incorporated during training to enhance model robustness. HGD [19] utilizes a denoising autoencoder guided by high-level representations to eliminate adversarial perturbations. Similarly, NRP [26], leverages a self-supervised adversarial training mechanism to purify input samples, demonstrating strong effectiveness against transfer-based attacks. Certified defense methods, such as RS [3], train robust classifiers by adding noise to the input and providing provable robustness within a certain radius. Furthermore, diffusion purification such as Diffure [27] employs diffusion models to process input samples and reduce adversarial perturbations, further enhancing model resilience against adversarial threats.
3 Methodology
In this section, we first introduce the prerequisite knowledge and our motivation. Then, we provide a detailed explanation of our CWT method. To highlight the differences between our approach and other methods, we also summarize the distinctions between our method and DIM [50], as well as BSR [39].
3.1 Preliminaries
Given a target model $f$ with parameters $\theta$ and a clean input $x$ with the ground-truth label $y$, the objective of an untargeted attacker is to generate an adversarial example $x^{adv}$ that is visually similar to $x$, such that $\|x^{adv}-x\|_{p}\leq\epsilon$ and causes the model to misclassify the input, meaning $f(x^{adv};\theta)\neq y$. Here, $\epsilon$ denotes the perturbation budget, and $\|\cdot\|_{p}$ is the $\ell_{p}$-norm that quantifies the allowed perturbation. In this paper, we focus on the $\ell_{\infty}$-norm (where $p=\infty$), which restricts the maximum change to each individual pixel.
To generate such an adversarial example, the attacker generally maximizes the loss function $J(x^{adv},y;\theta)$ (e.g., cross-entropy loss) with respect to $x^{adv}$, subject to the constraint $\|x^{adv}-x\|_{p}\leq\epsilon$. This can be formalized as follows:
$x^{adv}=\mathop{\mathrm{arg\,max}}_{\|x^{adv}-x\|_{p}\leq\epsilon}J(x^{adv},y;\theta)$ (1)
Many methods have been proposed to approximate this optimization. Among these, the Momentum Iterative Fast Gradient Sign Method (MI-FGSM)  [4] is particularly notable for enhancing the transferability of adversarial examples across diverse models. This method incorporates a momentum term into the gradient calculation to stabilize the perturbation direction across iterations, thereby increasing the probability of successful attacks on various models. The update rule for the accumulated gradient $g_{t}$ is:
$g_{t+1}=\mu\cdot g_{t}+\frac{\nabla_{x}J(x^{adv}_{t},y;\theta)}{\|\nabla_{x}J(x^{adv}_{t},y;\theta)\|_{1}}$ (2)
$x^{adv}_{t+1}=x^{adv}_{t}+\alpha\cdot\text{sign}(g_{t+1})$ (3)
where $\mu$ is the decay factor that controls the influence of previous gradients, and $g_{t}$ is the accumulated gradient at iteration $t$. The $\ell_{1}$-normalization of the gradient helps maintain consistency in the perturbation direction across iterations, enhancing the transferability of the attack.
3.2 Motivation
Despite differences in architectures and parameters, models trained for the same task often exhibit similar decision boundaries due to a shared latent space [16]. Adversarial attacks leveraging input transformations aim to generate diverse perturbations, enabling gradient-based methods to effectively explore this shared latent space and identify vulnerabilities in model decision boundaries.
However, experiments reveal a critical limitation: adversarial examples that perform well on the surrogate model often fail to generalize effectively to target models. This observation raises an important question:
Why does this discrepancy occur?
As illustrated in Figure 3, we observe that even when classification results are identical, the attention distributions of different models vary significantly. For instance, as shown in Figure 3 (a), ResNet-18 focuses on the belly of the snake, while Swin focuses on both the head and tail of the snake. Motivated by these findings, we hypothesize that the transferability of adversarial attacks improves when perturbations better cover the regions of interest across all models. This hypothesis aligns with the principle of leveraging gradients from multiple models to optimize interactive perturbations.
To investigate this hypothesis, we first evaluate the attention distributions of transformed images generated by various input-transformation-based attack methods, as shown in Figure 2. It reveals that methods such as DIM, SIM, Admix, Maskblock, and US-MM fail to significantly alter the attention distributions. While BSR enriches the attention regions, it often shifts the focus to irrelevant areas outside the objects of interest in the original image. These findings prompt a novel question:
How can we generate transformed images that encourage models to focus on diverse regions of the objects within an image?
3.3 Component-Wise Transformation
Encouraging models to focus on diverse regions of the object within an image is a challenging task for several reasons: 1) First, models tend to concentrate their attention on the most salient features of an object. Deep learning models are optimized to identify the most discriminative parts that contribute significantly to classification, often neglecting other informative regions. 2) Second, the attention distribution is highly dependent on model architecture and training data. Different models may attend to different regions of the same object, making it difficult to design a unified approach that effectively diversifies attention across various models.
To address these challenges, a straightforward method is to apply random masking to the image, forcing the model to utilize other regions for prediction. However, this approach leads to significant information loss, as substantial parts of the image are obscured, potentially impairing the attack algorithm performance.
Therefore, we define our objectives as:
  • Maximize the retention of original image information: Ensure that essential features and semantics of the original image are preserved.
  • Maximize the diversity of model attention regions: Generate multiple transformed images where the model’s attention covers as much as possible the region where the object is located in the original image.
Motivated by human perception [28], where people tend to focus on larger objects, previous work [55] has also shown that models tend to pay attention to larger regions within an image. This insight suggests that we could apply a zoom-in operation to the image to enlarge the object, potentially shifting the model’s attention. However, globally scaling up the image fails to effectively change the relative size differences between different parts of the object and can cause significant displacement of the object from its original position.
To overcome these limitations, we propose CWT, which applies a sequence of transformations to individual blocks of an image. As shown in Figure 1, these transformations include bilinear interpolation (for more details on interpolation, refer to Appendix B), comprising pre-interpolation (interpolation-based shrinking) and block-wise scaling (interpolation-based enlargement), as well as selective rotation.
Formally, the CWT process involves the following steps:
  1. 1.
    Image Partitioning: Divide the image $x$ into a grid of $n\times n$ non-overlapping blocks $\{B_{i,j}\}$, where $i,j\in\{1,2,\dots,n\}$.
  2. 2.
    Pre-Interpolation: For each block $B_{i,j}$, apply scaled down to reduce redundancy and compress features. Specifically, $B_{i,j}$ is scaled by a random factor $s_{i,j}$, where:
    $s_{i,j}\sim\text{Uniform}[s_{\text{min}},s_{\text{max}}],$ (4)
    The scaled-down block $B^{\prime}_{i,j}$ is calculated as:
    $B^{\prime}_{i,j}=\text{Interpolate}(B_{i,j},\text{size}=(H^{\prime}_{i,j},W^{\prime}_{i,j})),$ (5)
    where:
    $H^{\prime}_{i,j}=\lfloor H_{i,j}/s_{i,j}\rfloor,\quad W^{\prime}_{i,j}=\lfloor W_{i,j}/s_{i,j}\rfloor.$ (6)
    The effectiveness of the Pre-Interpolation step is demonstrated in Figure 5, which shows how it contributes to reducing redundancy and improving adversarial robustness.
  3. 3.
    Block-wise Scaling: After pre-interpolation, each block is scaled up using the same scaling factor $s_{i,j}$, focusing attention on specific features. The scaled-up block $\hat{B}_{i,j}$ is computed as:
    $\hat{B}_{i,j}=\text{Interpolate}(B^{\prime}_{i,j},\text{size}=(H^{\prime\prime}_{i,j},W^{\prime\prime}_{i,j})),$ (7)
    where:
    $H^{\prime\prime}_{i,j}=\lfloor H_{i,j}\cdot s_{i,j}\rfloor,\quad W^{\prime\prime}_{i,j}=\lfloor W_{i,j}\cdot s_{i,j}\rfloor.$ (8)
  4. 4.
    Selective Rotation: To further diversify the spatial representation of features and avoid the massive information loss caused by rotating all blocks, Selective Rotation is applied to a subset of blocks. Let $N$ be the total number of blocks, and let $\mathcal{R}\subseteq\{1,2,\dots,N\}$ denote the indices of $k$ randomly selected blocks for rotation ($|\mathcal{R}|=k$). For each selected block $B_{i,j}\in\mathcal{R}$, a random rotation angle $r_{i,j}$ is sampled from:
    $r_{i,j}\sim\text{Uniform}(-r_{\text{max}},r_{\text{max}}),$ (9)
    where $r_{\text{max}}$ defines the allowable rotation range. The block is rotated around its center, and any padding is filled with zeros to maintain the original dimensions.
  5. 5.
    Random Cropping: After scaling and optional rotation, each block $\hat{B}_{i,j}$ is cropped back to its original size $H_{i,j}\times W_{i,j}$.
  6. 6.
    Image Reconstruction: After all transformations are applied to each block, the blocks are reassembled into a single transformed image:
    $R_{k}(I)=\bigcup_{i=1}^{n}\bigcup_{j=1}^{n}\tilde{B}_{i,j}.$ (10)
To stabilize gradient calculations during adversarial training with CWT, we compute the average gradient over $N$ transformed images:
$\bar{g}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{x^{\text{adv}}}J(T(x^{\text{adv}},n,N,s_{\text{max}},k,r),y;\theta)$ (11)
Where $T(x^{\text{adv}},n,N,s_{\text{max}},k,r)$ represents the transformation applied to the augmented image, parameterized by the input image $x^{\text{adv}}$, the number of blocks $n$, the total number of transformed images $N$, the maximum scaling factor $s_{\text{max}}$ (with $s_{\text{min}}=1.0$), the maximum rotation angle $r$, and the number of rotated blocks $k$. $J$ denotes the loss function (cross-entropy), and $y$ is the ground truth label.
Here we integrate our CWT method into MI-FGSM, and summarize the algorithm in Algorithm 1.
Algorithm 1 Component-Wise Transformation
0:  A classifier $f$ with parameters $\theta$, loss function $J$, a raw example $x$ with ground-truth label $y$, the magnitude of perturbation $\epsilon$, learning rate $\alpha$, decay factor $\mu$, number of iterations $T$, number of transformed images $N$, number of blocks $n$, the maximum scaling factor $s_{\text{max}}$ , the maximum rotation angle $r$, and the number of rotated blocks $k$.
0:  An adversarial example $x^{\text{adv}}$.
1:  Initialize $\delta=0$, $g_{0}=0$, and $\alpha=\epsilon/T$
2:  for $t=1\to T$ do
3:     Calculate the gradient by Eq. 11.
4:     Update the momentum $g_{t}$ by:
$g_{t}=\mu\cdot g_{t-1}+\frac{\bar{g}}{\|\bar{g}\|_{1}}$
5:     Update the adversarial perturbation $\delta$ by:
$\delta=\alpha\cdot\text{sign}(g_{t})$
6:     Update the adversarial example by:
$x^{\text{adv}}_{t}=x^{\text{adv}}_{t-1}+\delta$
7:     Clip $\delta$ to ensure $\|\delta\|_{\infty}\leq\epsilon$:
$x^{\text{adv}}_{t}=\text{Clip}_{x,\epsilon}(x^{\text{adv}}_{t}-x,-\epsilon,\epsilon)$
8:  end for
9:  return  $x^{\text{adv}}_{T}$
3.4 Comparison of CWT with DIM and BSR
DIM enhances the transferability of adversarial attacks by applying resizing and padding to input samples, whereas BSR achieves this by randomly shuffling and rotating image blocks. Despite incorporating similar operations, our method, CWT, differs significantly from these approaches. Below, we highlight the distinctions between our method and the two methods.
  • CWT Vs. DIM In contrast to DIM, which performs scaling with fixed values on the entire image, our method emphasizes a random enlargement operation with pre-interpolation. Moreover, while DIM applies this operation to the entire image uniformly, our approach targets individual blocks within the image, introducing more localized transformations.
  • CWT Vs. BSR As for BSR, although both methods involve block partitioning, our approach discards the shuffle operation and primarily utilizes block interpolation to modify the image structure. Regarding rotation, instead of applying rotation to all blocks, our method adopts a selective rotation mechanism, where the number of rotated blocks is limited.
4 Experiments
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
RN-18 DIM 100* 61.7 66.1 90.4 30.4 37.4 53.4 56.9 62.0 23.8
SIM 100* 59.6 64.1 90.5 24.6 35.7 49.0 53.3 59.6 25.5
Admix 100* 69.9 74.6 95.4 31.2 42.6 59.8 63.0 67.1 23.6
MaskBlock 100* 48.8 50.4 79.6 18.6 25.3 38.4 43.7 50.6 27.1
US-MM 100.0* 66.6 71.7 94.3 29.1 41.0 56.8 60.3 65.0 24.2
BSR 100.0* 89.1 90.2 99.4 49.1 62.3 79.4 79.2 81.1 16.7
CWT 100.0* 90.2 93.7 99.4 55.9 68.8 84.1 83.6 84.5 14.3
RN-101 DIM 61.7 84.7* 63.0 65.5 30.2 40.9 48.0 48.2 55.3 16.9
SIM 62.5 91.7* 64.7 66.3 25.8 38.1 46.3 46.1 55.2 20.4
Admix 74.7 94.9* 77.5 77.4 36.0 49.7 60.9 58.6 66.2 18.6
MaskBlock 61.3 92.3* 56.6 61.9 19.5 29.8 36.5 38.5 49.6 23.2
US-MM 81.8 95.1* 80.1 83.7 35.3 50.7 62.7 60.6 68.8 19.9
BSR 86.7 94.6* 89.6 90.2 58.5 72.9 80.8 78.1 81.4 10.9
CWT 87.7 95.1* 91.0 91.6 67.3 78.5 85.5 81.9 84.8 8.3
RX-50 DIM 61.3 55.8 86.7* 63.0 26.0 35.2 45.7 44.4 52.3 18.8
SIM 59.5 57.8 94.0* 64.6 20.9 32.9 39.7 41.6 51.4 22.7
Admix 71.8 72.9 95.8* 75.6 29.6 44.4 53.7 53.9 62.2 20.8
MaskBlock 55.3 46.2 94.1* 54.5 16.4 24.2 32.5 33.1 44.5 24.4
US-MM 78.6 73.2 96.9* 82.1 31.5 46.8 57.6 55.6 65.3 21.2
BSR 85.7 86.0 96.5* 88.7 48.2 68.3 76.0 73.9 77.9 14.1
CWT 87.3 86.4 95.9* 90.2 57.7 72.3 80.9 78.4 81.1 11.2
DN-121 DIM 86.7 70.7 72.0 99.9* 34.0 43.6 60.6 57.2 65.6 21.6
SIM 90.0 70.3 72.8 100* 31.5 41.0 57.0 59.6 65.3 23.1
Admix 95.3 80.5 83.2 100* 39.0 51.6 68.7 68.7 73.4 20.8
MaskBlock 82.8 59.3 61.9 100* 23.8 34.4 47.9 49.3 57.4 24.8
US-MM 95.6 77.9 79.9 99.9* 38.1 48.8 67.6 65.3 71.6 21.3
BSR 98.0 89.7 92.2 100.0* 52.1 66.9 81.6 79.4 82.5 15.3
CWT 99.0 90.9 93.8 100.0* 56.4 70.8 85.3 82.2 84.8 14.0
Table 1: Attack success rates (%) on eight models under single model setting with various single input transformations. The surrogate models are cnn-based. * indicates the surrogate model.
4.1 Experimental Setup
Models. We evaluate our proposed CWT method across three categories of target models: 1) CNN-based models, including four widely recognized architectures: ResNet-18 (RN-18) [12], ResNet-101 (RN-101) [12], ResNeXt-50 (RX-50) [51], and DenseNet-121 (DN-121) [15]; 2) Transformer-based models, comprising ViT-B [5], PiT-B [13], Visformer (ViF-S) [2], and Swin-Tiny (Swin-T) [22]; 3) Defense models, which include four defense methods: AT [35], HGD [19], RS [3], and DiffPure [27]; All models are pre-trained on the ImageNet dataset and evaluated on single model.
Dataset. Following previous works [^1] https://github.com/Trustworthy-AI-Group/TransferAttack, We evaluate our proposed CWT on 1000 images belonging to 1000 categories from the validation set of ImageNet dataset. All images are classified correctly by the models.
Baselines. We compare CWT with other input transformation-based methods. Specifically, the image scale-based methods (DIM [50], SIM [20]), the mixed image-based methods (Admix [42], US-MM [40]), and the block-based methods (MaskBlock [7], BSR [39]). For fairness, all the input transformations are integrated into MI-FGSM [4].
Evaluation Settings. We set the maximum perturbation $\epsilon=16/255$, the number of iterations $epoch=10$, the step size $\alpha=\epsilon/epoch$, the batch size $batchsize=8$, and the decay factor $\mu=1$ for MI-FGSM. For our method, CWT generates 20 scaled copies per iteration, divides the image into 2x2 blocks, applies a scaling factor ranging from 1.0 to 1.3, and applies a maximum rotation angle of $26^{\circ}$, selectively rotating $k=2$ blocks. For other methods, we follow the parameters reported in the original papers.
4.2 Evaluations on CNN-based Models
We first evaluate the transferability of adversarial examples generated by various attacks with input transformations when the surrogate model is based on a CNN architecture. Specifically, we generate adversarial examples using RN-18, RN-101, RX-50, and DN-121 as surrogate models and evaluate their attack success rates (ASR) across eight target models. The results are summarized in Table 1, where the first column indicates the surrogate model, the second column lists the attack methods, and the remaining columns present the ASR of different classification models under attack and the last two columns are the mean ASR and standard deviation, respectively.
Our proposed method consistently achieves state-of-the-art performance, with the highest mean ASR and lowest standard deviation across almost all experimental settings. Notably, when the surrogate model is RN-18, our method achieves the highest ASR across all target models. Furthermore, when the target models are ViT-B and PiT-B, our method surpasses the previous state-of-the-art BSR more than 5% and outperforms other baseline methods by at least 17.4%. These results confirm the effectiveness of our approach in generating transferable adversarial examples when the surrogate model is CNN-based.
4.3 Evaluations on Transformer-based models
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
ViT-B DIM 56.9 47.8 49.6 59.7 89.5* 54.0 55.0 61.7 59.3 13.1
SIM 62.4 48.6 52.4 64.3 99.1* 57.8 58.7 71.9 64.4 15.7
Admix 65.6 52.4 55.3 65.9 98.9* 61.3 63.5 73.1 67.0 14.4
MaskBlock 59.3 41.4 44.1 59.1 98.9* 46.9 51.3 62.6 58.0 18.3
US-MM 69.2 54.7 57.6 68.7 97.4* 61.8 65.0 75.9 68.8 13.4
BSR 78.2 74.7 75.9 81.0 90.2* 81.5 79.2 81.6 80.3 4.4
CWT 78.3 75.3 77.2 81.6 93.1* 83.5 82.4 83.9 81.9 5.1
PiT-B DIM 59.5 50.5 54.5 62.2 47.9 91.6* 63.5 65.3 61.9 13.5
SIM 58.6 45.4 48.7 60.1 38.3 97.7* 56.8 60.9 58.3 17.8
Admix 60.4 47.7 51.8 60.4 42.9 94.6* 61.1 63.5 60.3 15.7
MaskBlock 59.2 40.4 45.2 57.3 35.4 99.2* 55.9 57.2 56.2 19.5
US-MM 66.1 54.1 56.4 64.8 45.0 93.9* 64.0 68.0 64.0 14.3
BSR 82.9 80.8 84.0 86.5 75.0 97.8* 90.2 90.4 86.0 6.5
CWT 85.8 85.1 87.3 90.7 85.3 97.8* 92.6 92.6 90.0 4.3
ViF-S DIM 71.2 63.7 67.1 75.2 53.4 71.0 95.1* 76.9 71.7 12.0
SIM 68.1 60.5 62.4 72.0 49.7 65.9 96.7* 74.8 68.8 13.7
Admix 75.1 67.0 70.6 78.3 56.5 72.8 97.0* 81.6 74.9 11.8
MaskBlock 64.2 47.3 51.3 64.4 35.7 54.9 99.2* 65.2 60.3 18.7
US-MM 82.1 70.5 73.6 83.3 56.2 73.7 97.6* 82.9 77.5 12.1
BSR 90.7 86.9 90.7 93.5 73.3 88.9 99.3* 92.7 89.5 7.0
CWT 89.1 87.6 89.9 94.0 80.4 91.9 99.1* 93.7 90.7 5.1
Swin-T DIM 67.2 53.4 56.8 68.8 48.7 65.3 69.3 96.2* 65.7 14.5
SIM 48.9 29.9 34.1 45.4 27.7 35.5 43.1 98.1* 45.3 22.6
Admix 55.1 33.9 37.4 50.0 28.3 37.8 47.2 98.2* 48.5 22.0
MaskBlock 47.8 26.5 30.9 44.5 24.4 31.6 39.1 98.3* 42.9 23.9
US-MM 57.5 32.2 37.8 52.2 26.8 37.8 47.4 96.8* 48.6 22.0
BSR 88.7 82.5 86.0 91.2 71.6 89.7 91.0 98.3* 87.4 7.3
CWT 90.7 85.9 88.1 93.6 80.8 92.9 94.1 98.5* 90.6 5.2
Table 2: Attack success rates (%) on eight models under single model setting with various single input transformations. The surrogate models are Transformer-based. * indicates the surrogate model.
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
DIM 36.4 32.9 32.0 34.6 33.4 33.2 33.5 33.7
SIM 36.5 31.6 32.1 35.6 35.4 33.4 34.3 31.0
ADMIX 38.4 32.6 32.2 38.0 35.9 33.5 34.3 31.8
MaskBlock 34.0 31.7 31.3 32.6 33.7 32.6 32.2 31.0
US-MM 38.3 34.9 33.7 37.9 36.4 34.8 35.4 32.0
BSR 40.3 36.9 35.2 38.9 38.0 35.1 36.7 37.5
CWT 41.1 38.0 36.9 39.9 37.7 36.9 37.7 39.4
Table 3: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under AT.
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
DIM 24.4 29.8 20.7 27.2 25.7 23.4 26.8 21.1
SIM 23.3 28.9 18.0 24.7 27.8 20.9 25.9 14.2
ADMIX 24.2 35.2 23.3 30.5 28.9 23.1 29.3 14.9
MaskBlock 18.8 22.7 15.5 20.2 23.5 21.8 21.1 12.9
US-MM 24.5 33.8 23.2 27.0 30.1 23.3 29.4 13.8
BSR 33.6 45.5 28.8 33.8 39.5 35.0 37.9 30.5
CWT 40.5 56.7 38.4 41.3 43.2 43.2 48.7 37.8
Table 4: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under Diffure. The classifier is RN-101.
Building on the results observed with CNN-based models, we next evaluate our method’s performance when the surrogate model is based on a transformer architecture. Specifically, we use Transformer-based models, including ViT-B, PiT-B, ViF-S, and Swin-T, as surrogate models to generate adversarial examples and evaluate their transferability. The results are summarized in Table 2.
From the results, we observe that our method also achieves SOTA performance in terms of both mean ASR and standard deviation across most of experimental settings. Notably, when the surrogate model is ViT-B, our method does not achieve the highest ASR on the original classification model (ViT-B itself) but consistently achieves the best performance on all other classification models. This demonstrates the robustness and generalizability of our approach, as it effectively transfers adversarial examples to different target architectures, even when the surrogate model has inherent challenges in transferability.
4.4 Evaluations on Defense Method
To comprehensively assess the robustness of our proposed method against diverse defense mechanisms, we conducted experiments under four defense settings: Adversarial Training (AT), High-level representation Guided Denoiser (HGD), Randomized Smoothing (RS), and Diffusion Purification (Diffure). More extensive evaluations can be found in Appendix A. As summarized in Tables 3 through 6, our method, particularly CWT, consistently achieves the highest attack success rates across most of defense settings, outperforming BSR and other baseline methods. For instance, under Diffure, it outperforms BSR by 8.1% in mean ASR and exceeding weaker baselines by more than 17%. Even under RS, a certified defense known for its robustness, our method maintains competitive ASRs (e.g., 28.7% for RN-101 and 29.5% for ViF-S), showcasing its resilience. These results validate the robustness and effectiveness of our approach across diverse defenses.
4.5 Ablation Study
To further evaluate CWT, we conduct ablation studies on five key hyperparameters and one critical step: number of blocks $n$, maximum scale factor $s_{\text{max}}$, maximum rotation angle $r$, number of rotated blocks $k$, number of transformed copies $N$, and the key step pre-interpolation. All experiments are conducted using RN-101 as the surrogate model.
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
DIM 56.9 49.7 42.1 66.7 43.0 45.7 57.2 46.2
SIM 52.7 43.5 36.4 64.3 42.4 37.2 49.1 22.5
ADMIX 63.1 58.2 50.6 76.5 46.8 39.9 57.2 25.9
MaskBlock 38.2 34.0 26.8 48.9 35.9 32.3 36.9 20.5
US-MM 59.4 60.1 52.2 72.5 48.2 44.8 61.2 24.5
BSR 86.9 80.9 74.9 90.1 71.9 74.2 81.0 75.0
CWT 91.2 85.4 81.9 93.1 73.9 81.8 85.0 83.5
Table 5: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under HGD.
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
DIM 26.0 21.9 22.0 25.2 23.6 22.2 22.9 22.9
SIM 26.0 21.3 21.3 26.0 25.2 22.7 24.0 21.5
ADMIX 27.9 22.7 22.5 27.8 26.0 22.6 24.2 21.1
MaskBlock 23.8 21.2 20.8 23.0 23.6 22.4 22.1 20.5
US-MM 27.6 23.8 23.7 26.4 26.8 24.1 25.8 21.8
BSR 27.8 26.2 25.1 26.7 27.3 26.0 26.1 24.9
CWT 30.9 28.7 27.9 30.8 29.3 27.8 29.5 28.7
Table 6: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under RS. The defense model used is ResNet-50 with a noise level of 0.50.
On the number of blocks $n$.
The impact of block partitioning is illustrated in Figure 4 (a). When $n=1$, CWT applies global operations to the entire image, which fails to sufficiently disrupt the attention heatmaps. Increasing $n$ to $2$ significantly improves the attack success rate , as the $2\times 2$ block division introduces localized transformations, encouraging the model to focus on diverse regions of the object.
Figure 4: Attack success rates (%) of various models on the adversarial examples generated by CWT with different numbers of blocks, maximum resize rates, maximum rotate angles, numbers of rotation blocks and numbers of transformed copies. The adversarial examples are crafted using the RN-101 model and tested on seven other models under the black-box setting.
When $n>2$, performance begins to degrade due to increased information loss from operating on a larger number of blocks. However, in contrast to BSR, the performance does not drop sharply, as our selective rotation method stabilizes the results. Consequently, we set $n=2$.
On the maximum scale factor $s_{\text{max}}$.
As shown in Figure 4 (b), the maximum scale factor significantly impacts the ASR. When $s_{\text{max}}>1.4$, the ASR starts to decline gradually. This decline is due to the larger resizing requiring a random cropping operation, and as the image is enlarged, the cropped areas become larger, leading to information loss. Therefore, we set $s_{\text{max}}=1.3$ as the optimal value for our experiments.
On the maximum rotation angle $r$.
As demonstrated in Figure 4 (c), we observe that gradually increasing $r$ improves ASR, but fluctuations occur when the angle exceeds 25°. Considering the experimental error margins, we interpret this fluctuation as insignificant, and thus we select $r=26^{\circ}$.
On the number of rotated blocks $k$.
As shown in Figure 4 (d), we find that increasing $k$ generally leads to a decline in performance on most target models. This is due to the additional padding introduced by more rotation operations, which results in increased information loss. Therefore, we select $k=2$ as the optimal value.
On the number of transformed copies $N$.
As shown in Figure 4 (e), the ASR steadily improves with increasing $N$, as it stabilizes gradient updates and introduces greater diversity. However, to maintain consistency with the BSR baseline and minimize computational overhead, we set $N=20$ in our experiments.
Figure 5: Attack success rates (%) of various models on the adversarial examples generated by CWT with and without Pre-interpolation. The surrogate model is RN-101.
On pre-interpolation.
As shown in Figure 5, we observe that pre-interpolation significantly improves the performance of the method. This demonstrates its ability to effectively remove redundant information, thereby enabling the model to focus on the relevant regions of the transformed image.
5 Conclusion
Based on the observation that inconsistent attention distributions across models are a major barrier to adversarial example transferability, we propose CWT, a method utilizing interpolation and selective rotation. CWT effectively generates diverse transformed images, enabling a single surrogate model to focus on different regions of the target object. Experimental results demonstrate that our approach not only surpasses existing SOTA methods in ASR but also achieves significantly lower standard deviation, highlighting its stability and generalization capabilities. Furthermore, CWT offers a novel perspective for input transformation-based adversarial attacks.
Supplementary Material
Appendix A Extended Evaluations on Defense Methods
In this section, we evaluate the performance of our proposed method on five adversarially trained models, namely Inc-v3ens, Inc-v3ens3, Inc-v3ens4, IncRes-v2ens and Res50in  [38, 9], as well as the NRP [26] method. The results are summarized in Tables  A1 through  A3.
Specifically, in Table  A1, we analyze the attack effectiveness of adversarial samples generated using different attack methods on various surrogate models against adversarially trained models. It is worth noting that these models do not achieve 100% classification accuracy on clean images. However, the misclassification rates of the models are all below 5%. The results indicate that our method surpasses previous SOTA approaches by an average improvement of 5.3%.
For Tables A2 and  A3, we apply NRP to process adversarial images generated by different defense methods and ASR on eight models. The results demonstrate that our method, CWT, outperforms existing techniques in terms of both ASR and standard deviation for CNN-based and Transformer-based surrogate models. Notably, while maintaining a reduced standard deviation, our approach consistently improves the ASR by an average of more than 5% over prior SOTA methods.
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
Inc-v3ens DIM 64.2 50.7 46.2 68.0 47.8 48.8 61.4 52.4 54.9 7.8
SIM 60.1 45.2 41.2 67.2 49.0 44.1 54.8 28.6 48.8 11.2
Admix 71.6 58.9 50.6 77.9 51.4 43.8 59.5 31.1 55.6 14.0
MaskBlock 44.4 36.3 31.1 51.0 42.5 39.2 41.8 25.9 39.0 7.4
US-MM 68.7 60.1 52.8 74.4 53.5 47.2 64.8 30.8 56.5 12.9
BSR 88.9 79.1 71.5 89.8 72.7 75.0 81.4 75.8 79.3 6.5
CWT 92.7 84.6 79.0 92.6 74.0 82.1 86.1 84.0 84.4 5.9
Inc-v3ens3 DIM 60.2 48.3 43.7 65.7 44.7 47.1 58.4 49.8 52.2 7.6
SIM 57.1 43.7 36.4 63.8 46.7 41.2 51.7 26.1 45.8 11.1
Admix 67.9 56.4 49.7 75.1 49.5 42.1 59.2 29.3 53.6 13.5
MaskBlock 43.4 36.2 28.5 51.1 36.5 38.4 40.9 24.8 37.5 7.7
US-MM 65.0 60.1 51.2 72.3 50.1 45.2 61.9 29.8 54.5 12.5
BSR 85.7 77.7 70.5 87.1 71.0 71.6 78.5 73.6 77.0 6.1
CWT 91.4 82.5 77.5 90.8 72.5 79.4 85.5 81.6 82.7 6.0
Inc-v3ens4 DIM 61.5 45.2 39.1 65.2 43.9 46.0 54.8 48.7 50.6 8.5
SIM 52.7 39.8 34.4 62.8 44.5 38.2 48.6 25.8 43.4 10.7
Admix 64.3 53.2 45.7 72.9 47.5 40.0 54.6 28.0 50.8 13.1
MaskBlock 39.1 32.3 27.8 47.0 36.8 33.7 39.3 24.9 35.1 6.6
US-MM 61.6 56.8 46.8 71.0 49.7 43.5 58.5 27.5 51.9 12.4
BSR 85.1 73.6 66.8 86.4 69.2 70.5 77.2 72.9 75.2 6.8
CWT 90.9 82.0 75.4 90.8 72.1 79.1 81.7 79.9 81.5 6.2
IncRes-v2ens DIM 48.0 43.9 34.8 55.1 41.5 39.8 50.2 39.8 44.1 6.2
SIM 42.1 36.7 30.3 50.6 39.7 33.6 44.6 19.2 37.1 9.0
Admix 50.6 50.1 41.6 61.8 44.9 34.2 50.6 21.0 44.4 11.6
MaskBlock 30.1 28.4 23.3 37.7 33.6 29.2 30.0 17.5 28.7 5.7
US-MM 49.3 51.7 44.0 59.1 44.4 38.4 53.2 21.0 45.1 10.9
BSR 75.0 71.4 63.9 77.8 67.4 66.1 69.5 64.0 69.4 4.8
CWT 81.8 79.6 73.5 84.6 69.6 73.8 76.9 73.9 76.7 4.7
Res50in DIM 83.8 54.1 51.1 79.6 49.7 52.9 61.8 57.6 61.3 12.3
SIM 81.6 53.9 49.3 79.1 54.9 46.8 57.6 36.2 57.4 14.6
Admix 90.1 67.1 61.3 87.7 58.2 49.2 64.6 40.3 64.8 16.1
MaskBlock 72.4 47.1 39.8 70.5 48.3 44.7 47.6 33.3 50.5 13.0
US-MM 88.2 73.1 66.8 87.7 60.7 52.6 70.7 41.4 67.7 15.1
BSR 97.1 83.2 78.8 95.3 75.1 77.0 82.3 82.3 83.9 7.6
CWT 98.1 85.7 82.2 95.3 74.6 80.5 85.7 85.9 86.0 7.1
Table A1: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under adversarially trained models.
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
RN-18 DIM 98.9* 39.5 40.7 64.5 19.5 25.8 30.4 34.7 44.3 25.8
SIM 99.5* 40.6 44.7 67.5 17.2 24.9 31.9 37.3 45.5 26.5
Admix 99.9* 47.8 51.9 77.7 19.9 28.4 35.8 41.5 50.4 26.5
MaskBlock 98.0* 32 35.5 57.5 13.6 19 25 28.1 38.6 27.4
US-MM 99.4* 46.6 51.3 76.2 21.5 29.3 39.5 42.6 50.8 25.5
BSR 98.8* 53.4 60.0 81.8 25.9 37.7 46.7 50.0 56.8 23.5
CWT 99.4* 58.6 64.9 87.4 33.5 43.7 52.4 55.2 61.9 21.9
RN-101 DIM 54.4 56.8* 38.7 49.5 19.1 25 28.5 29.4 37.7 14.4
SIM 55.7 62.6* 40.8 51 17.1 25.7 28.3 30.6 39.0 16.1
Admix 60.9 75.7* 52.9 58.9 22.6 33.2 37.9 39.8 47.7 17.4
MaskBlock 53.3 54.6* 35.6 47.3 13 19.2 23 24.9 33.9 16.2
US-MM 99.4* 46.6 51.3 76.2 21.5 29.3 39.5 42.6 50.8 25.5
BSR 71.9 73.4* 63.6 70.8 32.2 45.0 51.8 50.1 57.7 14.9
CWT 74.7 80.9* 68.9 75.3 40.4 52.4 58.2 57.3 63.5 13.7
RX-50 DIM 54 32.8 57.4* 46.8 16 23 26.8 27.2 35.5 15.3
SIM 55.9 35.5 63.3* 48.3 14.9 22.6 25.5 28.1 36.8 17.3
Admix 60.1 45.4 76.2* 56.6 19.8 29.7 34.1 35.9 44.7 18.5
MaskBlock 53.4 29.3 55.2* 41.9 11.9 18.4 22.6 23.7 32.1 16.2
US-MM 68.9 49.0 78.8* 63.0 31.3 37.1 39.9 48.6 48.7 20.0
BSR 68.8 54.1 74.3* 67.2 28.3 40.8 46.3 46.4 53.3 15.8
CWT 70.8 59.8 80.4* 71.5 34.3 47.0 53.7 52.6 58.8 15.0
DN-121 DIM 68.4 41.2 45.3 97* 20.2 26.8 33.9 35.7 46.1 25.1
SIM 74.8 45.5 49.6 98.6* 20.9 30.1 38.4 40.6 49.8 25.2
Admix 81.3 55.9 57.3 99.3* 28 35 43.9 48.1 56.1 23.8
MaskBlock 67.5 35.3 39.1 95.1* 15 22.7 29.3 31.7 42.0 26.4
US-MM 82.0 54.4 58.6 98.7* 27.5 35.2 45.8 47.8 56.4 23.8
BSR 83.8 56.3 61.0 96.6* 29.9 38.6 48.5 48.7 57.9 22.4
CWT 83.1 60.9 64.7 98.6* 35.1 46.2 52.9 54.8 62.0 20.4
Table A2: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under NRP. The surrogate models are CNN-based. * indicates the surrogate model.
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
ViT-B DIM 52.5 34.3 34.1 47.9 70.9* 34.4 35.4 39 43.6 13.1
SIM 57.6 36.1 39.2 52.5 85.3* 38.4 40.5 47 49.6 16.3
Admix 57.9 37.9 41.4 54.2 86.9* 40.7 41.8 51.3 51.5 16.0
MaskBlock 56.8 30.3 32.7 47.2 81.6* 30.1 32.6 40.6 44.0 17.9
US-MM 61.1 40.0 42.8 58.6 84.6* 42.5 44.7 55.0 53.7 15.5
BSR 66.0 49.9 51.9 62.8 69.1* 55.3 55.1 56.8 58.4 6.9
CWT 66.7 52.1 53.8 64.1 75.9* 59.1 57.7 58.8 61.0 7.7
PiT-B DIM 56 32.2 36.1 48.1 27.2 67.8* 37.2 38.1 42.8 13.5
SIM 53.1 32 34.5 45.8 24.7 73.5* 35.9 38.1 42.2 15.3
Admix 54.8 34.7 36 47 25.8 73.8* 39.8 42.8 44.3 14.7
MaskBlock 56.8 29.7 34.2 44.5 23.1 76.5* 34.8 37.1 42.1 17.2
US-MM 68.4 37.4 39.2 52.2 29.4 74.4* 43.6 46.0 47.6 14.0
BSR 68.1 50.2 54.9 65.1 44.1 79.7* 60.0 61.9 60.5 11.1
CWT 70.6 56.7 59.9 69.3 52.7 85.1* 65.8 67.3 65.9 10.0
ViF-S DIM 59.3 37.8 41.5 53.6 28.6 41.1 76.7* 44.9 47.9 14.9
SIM 60.7 38.2 40.3 53 27.6 41.2 80.8* 47.8 48.7 16.4
Admix 61.8 42.5 46.3 59.3 32 47.2 85.6* 54.1 53.6 16.1
MaskBlock 57.9 32.8 33.8 47.6 20.1 30.8 76.3* 38.5 42.2 17.8
US-MM 68.8 46.5 52.7 64.1 35.3 50.9 86.4* 58.7 57.9 15.5
BSR 71.7 52.1 56.8 69.2 40.5 59.3 81.6* 61.4 61.6 12.7
CWT 73.4 59.0 63.0 73.1 50.8 65.9 87.9* 68.9 67.8 11.1
Swin-T DIM 54.6 31.5 33.2 46.1 24.6 33.5 37.4 74.5* 41.9 16.1
SIM 45.7 25.4 25.9 37.4 14 22 24.5 70.4* 33.2 17.9
Admix 50.9 25 26.7 40.5 15.2 23.8 27.7 75.2* 35.6 19.4
MaskBlock 50.1 24.1 24.9 38.3 13.5 20.8 24.9 69.1* 33.2 18.4
US-MM 52.1 26.1 29.6 43.1 16.2 23.2 28.3 74.3* 36.6 19.0
BSR 68.8 48.6 51.5 64.6 36.7 54.7 57.0 76.9* 57.4 12.6
CWT 70.9 54.6 57.7 71.5 45.1 62.4 62.6 83.1* 63.5 11.7
Table A3: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under NRP. The surrogate models are Transformer-based. * indicates the surrogate model.
Appendix B More Analysis on Interpolation
To better understand the role of interpolation in affecting model attention distributions, we conducted experiments using various interpolation methods (e.g., bilinear, bicubic, nearest neighbor, and area interpolation) under different scaling factors. As shown in Figure  A1, scaling factors below 1.0 tend to disperse attention across irrelevant regions, reducing focus on critical object areas. Conversely, scaling factors significantly above 1.0 excessively concentrate attention on limited regions. Moderate scaling factors, typically ranging from 1.0 to 1.8, yield the most balanced attention distributions, effectively redistributing focus across diverse object regions. For example, as depicted in Figure  A1 (a), a scaling factor of 1.0 primarily directs attention to a person’s back, while a factor of 1.4 shifts focus towards the hips, and 1.8 further moves attention to the left shoulder.
In our method, we strategically apply a two-step process to achieve balanced attention redistribution. First, we perform a shrinking operation on each block to disperse attention and eliminate redundant information. Next, we enlarge the block, refocusing the previously dispersed attention on critical regions of the object. This two-step process effectively balances attention distribution and enhances the model’s robustness.
Moreover, the choice of interpolation method plays a critical role in shaping attention distributions. Bicubic interpolation, while smooth, often over-smooths attention maps, making it less effective at capturing distinct object regions. Nearest neighbor interpolation demonstrates insensitivity to scaling factors above 1.0. Area interpolation, on the other hand, is overly sensitive to scaling factors below 1.0, resulting in attention maps that collapse onto irrelevant regions and fail to preserve essential object features. In contrast, bilinear interpolation achieves a balance between smoothness and precision, producing the most consistent and well-distributed attention maps.
Thus, we adopt bilinear interpolation in our method to ensure optimal attention distribution and robust adversarial performance.
Figure A1: Heatmaps of Different Interpolation Rates Using ResNet-18. (a) Area interpolation; (b) Bicubic interpolation; (c) Bilinear interpolation; (d) Nearest-neighbor interpolation.
Appendix C Examples of Adversarial Images
In this section, we visualize the adversarial images generated by our method, CWT, on eight surrogate models, using heatmaps on ResNet-18, as shown in Figure A2. It can be observed that CWT effectively transfers the target attention to different regions of the image.
Figure A2: Heatmaps of the adversarial images generated by CWT on eight surrogate models. The heatmaps generated on ResNet-18.